A. An airliner must reach a speed of 110 m/s to take off. If the available runway is 2.4 km



and the aircraft accelerates uniformly from rest at one end, what minimum acceleration



must be available if it is to take off?



B. A speeding motorist driving a car, passes a traffic police officer on a stationary



motorcycle. The police officer immediately gives chase: his uniform acceleration is



4ms-2



, and by the time he draws level with the speeding car, the police officer is



travelling at 30ms-1



.



i. How long did it take for the police officer to catch the car?



ii. If the speeding motorist continued to travel at a steady speed during the



chase, what was his speed?



C. A stone is thrown vertically upwards with a speed of 24.0 ms-1



. Using g = 9.8 ms-2



and



ignoring air resistance:



i. How much time is required to reach this height?



ii. Why are there two answers to (ii.)?



iii. Sketch the velocity-time graph for the stone, from it is released until it



returns to the thrower.

1
Expert's answer
2022-02-11T15:33:31-0500

A.

"a=\frac{v^2}{2l}=2.52~\frac m{s^2}."

B.

1.

"t=\frac va=7.5~\frac ms,"

2.

"v_1=v=30~\frac ms."

C.

"t=\sqrt{\frac {2h}g}=2.2~s."